ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: 34, 93 Aisagaliev, Serikbai, Kazakh State University, Kazakhstan: Theory and constructive methods of solving the boundary problems of controllability processes. The problems of controllability, time-optimal control, optimal control and classic boundary problems (two-pointed, periodical) for processes described by ordinary differential equations are researched. Theory and methods of solving two central problems of boundary problems: the necessary and sufficient conditions of solvability and construction the solution as a limit of minimised sequences are developed. Reference: S. A. Aisagaliev, Optimal control of linear systems with fixed ends of trajectory and restricted control, Differential equations, 32 (1996), 1–8. S. A. Aisagaliev, Controllability of some system of differential equations, Differential equations, 27 (1991), 1475–1486. S. A. Aisagaliev, Boundary problems of ordinary differential equations, Reports of NAS of RK, 2 (1993), 3–8.